# Training PINNs with Hard Constraints and Adaptive Weights: An Ablation Study

### v2

@@ Abstract
Training Physics-Informed Neural Networks (PINNs) to solve stiff time-dependent partial differential equations (PDEs) remains a significant challenge. The main difficulties lie in precisely enforcing initial conditions and balancing the various loss components. In stiff PDEs, where solutions show rapid transitions or sharp gradients, this balance becomes especially difficult. The wide range of scales in the dynamics can cause certain losses to dominate, leading to unstable or inefficient training. This paper presents a comprehensive ablation study focused on two pivotal training schemes: the enforcement of hard constraints for initial and boundary conditions, and the implementation of adaptive weights. We specifically examine their impact on the performance of PINNs when applied to stiff time-dependent PDEs from materials science and mathematical biology applications. We conduct extensive numerical experiments across a diverse range of time-dependent PDEs from Allen-Cahn, Cahn-Hillard, to Gray-Scott systems. We further discuss the implications of our findings for improving the robustness and efficiency of PINN training, particularly in settings where accurate representation of initial conditions and balanced loss contributions are paramount.
@@ 1 Introduction
Time-dependent partial differential equations (PDEs) constitute the mathematical foundation of numerous scientific and engineering disciplines, serving as essential tools for modeling a diverse range of physical phenomena. Their applications span from the intricate dynamics of fluid flow, as described by the Navier–Stokes equations, to the complex interplay of chemical reactions in biological systems, captured by reaction–diffusion models. Time-dependent PDEs offer a powerful framework for understanding and predicting the evolution of systems over time whitham2011linear ; burgers1948mathematical ; allen1979microscopic ; shen2010numerical ; bazant2017thermodynamic ; horstmann2013precipitation ; tian2015electrochemically ; hyman2014liquid ; takatori2015towards ; cahn1958free ; lee2014physical ; miranville2017cahn ; kudryashov1990exact ; michelson1986steady ; de1991turing ; nishiura1999skeleton ; zhabotinsky2007belousov ; cassani2021belousov ; kato1987nonlinear ; kevrekidis2001discrete ; fibich2015nonlinear . In particular, the imposition of periodic boundary conditions allows for the simulation of phenomena over effectively infinite physical domains, providing valuable insights into the intrinsic behavior of physical systems. However, even the numerical solution of these PDEs, especially in scenarios involving stiffness or complex non-linearities, can be exceedingly challenging, often demanding significant computational resources and specialized numerical techniques.
@@
The advent of machine learning has ushered in a new era of scientific computing, offering powerful tools for data-driven modeling and analysis. Machine learning methodologies have shown remarkable success in various science/engineering fields, revolutionizing areas such as protein folding prediction PF2019 , drug discovery and development CARRACEDOREBOREDO20214538 , and even optimizing fundamental algorithmic processes such as matrix-vector multiplication MMML2022 . However, traditional machine learning approaches, which rely on vast datasets, can prove impractical in scientific contexts where data acquisition is expensive, time consuming, or simply infeasible. Furthermore, purely data-driven models often lack physical fidelity and model interpretability, as they may fail to capture the underlying physical principles governing the systems they aim to describe. Physics-informed machine learning (PIML), specifically physics-informed neural networks (PINNs), address these limitations by seamlessly integrating physical laws into the learning process, significantly reducing the amount of training data needed. By embedding constraints derived from the governing PDEs directly into the neural network architecture, PINNs enable accurate predictions even with sparse and noisy data yang2019physics ; chen2021solving ; raissi2018hidden ; raissi2019physics ; liu2019multi ; liu2021dual ; jin2021nsfnets ; cai2021deepm ; rad2020theory ; zhu2021machine ; chen2022using ; wang2023pinnsbased . Despite the promising applications, effective training of PINNs, especially for stiff time-dependent PDEs, remains a significant challenge. The inherent difficulty in propagating information from initial and boundary data points into the computational domain, coupled with the need to balance multiple loss terms, often leads to convergence issues and suboptimal solutions wang2020understanding ; wang2020and ; mcclenny2020self ; coutinho2022physics .
@@
This paper presents an ablation study of training algorithms for PINNs applied to stiff partial differential equations (PDEs), focusing specifically on two schemes: hard constraints and self-adaptive loss weights mcclenny2020self . The hard-constraint scheme directly embeds information about initial and boundary conditions into the neural network architecture. This transformation reduces the effects of spectral bias from training various kinds of loss terms, leading to improved solution accuracy wang2020and . PINN’s training can treated as a regularized regression, where initial and boundary conditions, along with observational data, serve as training data, and the physics-informed loss acts as regularization. However, the multi-objective nature of the loss function and the sensitivity of PDE solutions to initial conditions can significantly hinder the search for global minimizers. By explicitly incorporating boundary and initial condition data into the network transformation, we streamline the training process, particularly for stiff time-dependent PDEs, where accurately capturing the initial conditions is critical. In addition, we investigate the use of adaptive loss weights to alleviate the spectral bias from the multi-objective loss in PINN training mcclenny2020self . Through our ablation study of these training schemes, we arrive at the conclusion that hard constrain transformation provides better improvement. We emphasize that hard constraints and self-adaptive weights are not the only effective strategies for handling stiff PDEs. These methods can be integrated with other advanced PINN techniques, such as causal PINNs wang2020and , time-marching PINNs wight2020solving , residual-based-attention (RBA) PINNs ANAGNOSTOPOULOS2024116805 , curriculum training krishnapriyan2021characterizing , and co-training PINNs zhong2024label , to achieve synergistic improvements in prediction accuracy and robustness. Through a rigorous ablation study, we demonstrate the efficacy of the proposed approach across a range of stiff time-dependent PDEs, offering valuable insights into optimal training strategies for PINNs in these challenging scenarios.
@@ 1.1 Related Work
PINNs have shown significant potential to solve a wide range of PDEs. However, challenges arise, particularly when dealing with “stiff” PDEs, where explicit numerical methods require extremely small time steps for stability. A primary concern is the imbalance between the fitting for initial data and residual components within the PINN loss function wight2020solving ; shin2020convergence ; wang2020understanding ; wang2020and . This imbalance often leads to a gradient descent that favors minimization of residual loss over data fitting loss, hindering convergence to accurate solutions. This issue is exacerbated in forward problems, where data is primarily concentrated at initial and boundary conditions, leaving limited information within the PDE domain. Specifically, in time-dependent problems, PINNs struggle to propagate information from initial conditions to subsequent time steps, a phenomenon widely documented in the literature wight2020solving ; mcclenny2020self ; krishnapriyan2021characterizing ; wang2022respecting ; haitsiukevich2022improved .
@@
Several strategies have been proposed to address these challenges. One approach, introduced by Wight and Zhao wight2020solving , involves time-marching, where the time domain is divided into smaller intervals, and PINNs are sequentially trained on each interval, starting from the initial condition. While effective, this method can be computationally expensive due to the need to train multiple networks. To improve information propagation, researchers have emphasized the importance of weighting initial condition data and residual points near the initial time wight2020solving ; mcclenny2020self . McClenny and Konstantinidis mcclenny2020self demonstrated that their self-adaptive PINNs can autonomously learn these weights during training, particularly in scenarios with complex initial conditions, such as the wave equation.
Further advancements include the “sequence-to-sequence” approach proposed by Krishnapriyan et al. krishnapriyan2021characterizing , which sequentially learns each time step in a time-marching scheme. Wang et al. wang2022respecting identified the lack of “causality” in standard PINN training and proposed a corrective weighting scheme that gradually prioritizes data from early to later times, leveraging information from the PDE residue. This approach shares similarities with the self-adaptive weighting mechanism of McClenny and Konstantinidis mcclenny2020self . Additionally, Haitsiukevich et al. haitsiukevich2022improved introduced a pseudo-label approach, resembling self-training, to further enhance PINN convergence and accuracy. In zhong2024label , a mixed training scheme using labal-propagation between data points and collocation points is introduced and tested. For stiff parabolic PDEs such as viscous Burgers with small viscosity, the authors in coutinho2022physics proposed an data-driven artificial viscosity scheme to produce stable solutions and accurate to capture sharp transitions in the solutions.
@@
Beyond loss weighting and time marching, structural modifications to PINNs have also been explored. Braga-Neto et al. braganeto2023characteristicsinformed introduced Characteristic-Informed Neural Networks (CINNs) for transport equations, where PDE information is integrated into the neural network architecture. A similar approach, albeit for Dirichlet boundary conditions, was proposed by Lagaris et al. Lagaris_1998 . Our method builds upon these architectural modifications, with a focus on enhancing stability and ensuring proper propagation of information, especially for stiff time-dependent PDEs. We carefully design our architecture to address the challenges of information propagation and loss balancing, aiming to improve the robustness and efficiency of PINN training.
@@ 2 Methodology
We consider the following setup for a family of time dependent PDEs. Let $u$ be an unknown function defined on $[0,T]\times\Omega$ with $\Omega\subset\mathcal{R}^{d}$. Here the physical domain $\Omega$ comes with a Lipschitz boundary $\partial\Omega$ and $\bar{\Omega}=\Omega\cup\partial\Omega$. Let $\mathcal{P}$ be an operator which can include differentiation and integration. For example, for the $1D$ heat equation $u_{t}-\lambda u_{xx}=0$, $\mathcal{P}=\lambda\frac{\partial^{2}}{\partial x^{2}}$. Notice that $\mathcal{P}$ can be nonlinear. Let $\mathcal{B}$ be another operator on the boundary, that is, for Dirichlet boundary condition $u(t,x)=g(t,x)$ for $(t,x)\in[0,T]\times\partial\Omega$, $\mathcal{B}=\mathcal{I}$ (the identity). Then we say that $u$ is a solution of an PDE, if $u$ satisfies the following
$\begin{cases}u_{t}-\mathcal{P}(u)(t,\mathbf{x})&=f(t,\mathbf{x}),\quad(t,\mathbf{x})\in(0,T]\times\Omega,\\ u(0,\mathbf{x})&=u_{0}(\mathbf{x}),\quad\mathbf{x}\in\bar{\Omega},\\ \mathcal{B}(u)(t,\mathbf{x})&=g(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega,\\ \end{cases}$ (1)
Here the functions $u_{0}$, $f$, and $g$ are user-inputs and are assumed to satisfy desired regularity so that the existence and uniqueness of solutions for such a PDE is guaranteed. We also assume that the compactibility condition, that is, $g(0,\mathbf{x})=u_{0}(\mathbf{x})$ for $\mathbf{x}\in\partial\Omega$, is satisfied to prevent any ill-conditioning of the PDE problem.
@@
$\mathcal{B}(u)(t,\mathbf{x})$ $=u(t,\mathbf{x}),\quad\text{Dirichlet type},$
$\mathcal{B}(u)(t,\mathbf{x})$ $=\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Neumann type, $\mathbf{n}$ is the outward normal vector to $\partial\Omega$},$
$\mathcal{B}(u)(t,\mathbf{x};a,b)$ $=au(t,\mathbf{x})+b\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Robin type}.$
The periodic boundary condition gives $d$ different equations, and they are given as follows,
$\mathcal{B}(u)(t,\mathbf{x})=u(t,\mathbf{x})-u(t,\mathbf{x}+P\mathbf{e}_{i})=0,\quad i=1,\cdots,d.$ (2)
Here $P$ is the period and $\mathbf{e}_{i}$ is the $i^{th}$ standard basis vector in $\mathcal{R}^{d}$, that is,
@@
One can even consider a mixture of two or all of the above BC conditions. For example, we can have $\Omega=\Gamma_{1}\cup\cdots\cup\Gamma_{K}$ with each $\Gamma_{i}$ defines a different kind of BC. In this paper, we will mainly focus on the periodic type of boundary conditions, however other kinds of BCs can be
Classical numerical techniques such as the Finite Difference Method (which approximates derivatives on a point mesh), Finite Element Method (utilizing basis approximation on a triangular mesh with weak formulation), and Spectral Method (employing Fourier basis with FFT on a point mesh) are widely utilized in scientific computations. However, recent advancements in scientific machine learning integrate the principles of physics, particularly the understanding of PDEs, into the training process of machine learning models. This integration enables the development of PINNs to solve for $u$ as follows: one tries to find an approximate solution from a set of deep neural networks (of the same depth, same number of neurons on each hidden layer, and same activation functions on each layer) $\mathcal{H}_{\text{NN}}$, which is a minimizer of the following loss functional
$\text{Loss}(u_{nn})=\text{Data Loss}+\lambda*\text{Physical Loss}$
@@
Data $\text{Data Loss}=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
$\quad+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2}$
Physical $\text{Physical Loss}=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}$
for $u_{nn}\in\mathcal{H}_{\text{NN}}$. Here $\{(t^{CL},\mathbf{x}^{CL})_{i}\}_{i=1}^{N_{CL}}\in(0,T]\times\Omega$ are called collocation points, $\{(0,\mathbf{x}^{IC})_{i}\}_{i=1}^{N_{IC}}\in\{0\}\times\bar{\Omega}$ are the initial condition points, and $\{(t^{BC},\mathbf{x}^{BC})_{i}\}_{i=1}^{N_{BC}}\in[0,T]\times\partial\Omega$ are the boundary condition points. The minimizer, denoted as $u_{NN}$, will an be approximate solution to (1).
2.1 Hard Constrains
The main motivation for us to consider hard constrains transformation for PINNs is the possible ill-conditioning of the PDE solution map’s dependence on IC and BC. Traditional numerical methods do not need to encounter this kind of difficulty, as the IC data is exactly satisfied in their solutions. PINNs, on the other hand, use $L^{2}$-loss to fit the IC/BC and PDE residual data, which causes the solutions to be highly sensitive to the tightness of fit for the IC in the training (as pointed out in mcclenny2020self ). By transforming the need of fitting IC/BC conditions out of PINNs, it reduces the spectral bias in the training. Therefore, We consider the following transformation
$\tilde{u}(t,\mathbf{x})=\psi(t,\mathbf{x})+\phi(t,\mathbf{x})u_{nn}(t,\mathbf{x})$

### v3

# Stability in Training PINNs for Stiff PDEs: Why Initial Conditions Matter

@@ Abstract
Training Physics-Informed Neural Networks (PINNs) on stiff time-dependent PDEs remains highly unstable. Through rigorous ablation studies, we identify a surprisingly critical factor: the enforcement of initial conditions. We present the first systematic ablation of two core strategies, hard initial-condition constraints and adaptive loss weighting. Across challenging benchmarks (sharp transitions, higher-order derivatives, coupled systems, and high frequency modes), we find that exact enforcement of initial conditions (ICs) is not optional but essential. Our study demonstrates that stability and efficiency in PINN training fundamentally depend on ICs, paving the way toward more reliable PINN solvers in stiff regimes.
@@ 1 Introduction
Time-dependent partial differential equations (PDEs) constitute the mathematical foundation of numerous scientific and engineering disciplines, serving as essential tools for modeling a diverse range of physical phenomena. Their applications span from the intricate dynamics of fluid flow, as described by the Navier–Stokes equations, to the complex interplay of chemical reactions in biological systems, captured by reaction–diffusion models (Whitham, 2011; Burgers, 1948; Allen & Cahn, 1979; Shen & Yang, 2010; Bazant, 2017; Horstmann et al., 2013; Tian et al., 2015; Hyman et al., 2014; Takatori & Brady, 2015; Cahn & Hilliard, 1958; Lee et al., 2014; Miranville, 2017; Kudryashov, 1990; Michelson, 1986; De Kepper et al., 1991; Nishiura & Ueyama, 1999; Zhabotinsky, 2007; Cassani et al., 2021; Kato, 1987; Kevrekidis et al., 2001; Fibich, 2015). However, solution of stiff time-dependent PDEs can be exceedingly challenging to obtain and often requires substantial computational resources and specialized numerical techniques.
@@
Machine learning methodologies have demonstrated remarkable success across various scientific and engineering fields, transforming areas such as protein folding prediction (Jumper et al., 2021), drug discovery and development (Carracedo-Reboredo et al., 2021), and even the optimization of fundamental algorithmic processes such as matrix-vector multiplication (Fawzi et al., 2022). However, traditional machine learning approaches, which rely on large datasets, can be impractical in scientific contexts where data acquisition is expensive, time consuming, or simply infeasible. Moreover, purely data-driven models often lack physical fidelity and interpretability, as they may fail to capture the underlying physical principles governing the systems they aim to describe. Physics-informed machine learning (PIML), and in particular physics-informed neural networks (PINNs), address these limitations by integrating physical laws directly into the learning process. This significantly reduces the amount of training data required. By embedding constraints derived from governing PDEs into the neural network architecture, PINNs enable accurate predictions even when data are sparse or noisy (Yang et al., 2019; Chen et al., 2021; Raissi & Karniadakis, 2018; Raissi et al., 2019; Liu & Wang, 2019, 2021; Jin et al., 2021; Cai et al., 2021; Rad et al., 2020; Zhu et al., 2021; Chen et al., 2022; Wang et al., 2023). Despite their promise, effectively training PINNs for stiff time-dependent PDEs remains a major challenge. The difficulty of propagating information from initial and boundary conditions into the computational domain, combined with the need to balance multiple loss terms, often results in convergence issues and suboptimal solutions (Wang et al., 2020a, b; McClenny & Braga-Neto, 2020; Coutinho et al., 2023).
@@
This paper presents a complete study of investigating the importance of the training components for PINNs on solving stiff time-dependent PDES. First, we conduct an ablation study of training algorithms for PINNs applied to stiff time-dependent PDEs, focusing specifically on two schemes: hard constraints and self-adaptive loss weights (McClenny & Braga-Neto, 2020). The hard-constraint scheme directly embeds information about initial and boundary conditions into the neural network architecture. This transformation reduces the effects of spectral bias from training a multi-objective loss, leading to improved solution accuracy (Wang et al., 2020b). In addition, we investigate the use of adaptive loss weights to alleviate spectral bias in PINN training (McClenny & Braga-Neto, 2020). Through our ablation study of these training schemes, we conclude that the hard-constraint transformation provides the best improvement. Next, we provide a Neural Tangent Kernel perspective on how the hard-constrained transformation is reducing the spectral bias in PINN training. We emphasize that hard constraints is not the only effective strategies for handling stiff PDEs. Such method can be integrated with other advanced PINN techniques, such as causal PINNs (Wang et al., 2020b), time-marching PINNs (Wight & Zhao, 2020), residual-based-attention (RBA) PINNs (Anagnostopoulos et al., 2024), curriculum training (Krishnapriyan et al., 2021), and co-training PINNs (Zhong et al., 2024), to achieve synergistic improvements in prediction accuracy and robustness.
The rest of the paper is structured as follows. We compare our methodology to other related works in section 1.1; in section 2, we discuss the hard-constrained framework and other training methodologies used in our ablation and comparison study; we presented a detailed discussion on the theoretical mechanism behind the hard-constrained transformation in section 3.1; we presented the results (in summary) from four major categories of examples in section 4; we conclude our paper in 5, pointing out several future directions for this line of research. Since the examples we conducted are too extensive, the additional details are presented in Appendix section B.
Figure 1: 77 Benchmark Results: Truth vs HC-PINN vs Absolute Error.
@@ 1.1 Related Work
PINNs have shown significant potential to solve a wide range of PDEs. However, challenges arise, particularly when dealing with “stiff” PDEs, where explicit numerical methods require extremely small time steps for stability. A primary concern is the imbalance between the fitting of initial data and residual components within the PINN loss function (Wight & Zhao, 2020; Shin et al., 2020; Wang et al., 2020a, b). This imbalance often leads gradient descent to prioritize minimizing the residual loss over data fitting loss, hindering convergence to accurate solutions. This issue is exacerbated in forward problems, where data is primarily concentrated at initial and boundary conditions, leaving limited information within the PDE domain. Specifically, for time-dependent problems, PINNs struggle to propagate information from initial conditions to subsequent time steps, a phenomenon widely documented in the literature (Wight & Zhao, 2020; McClenny & Braga-Neto, 2020; Krishnapriyan et al., 2021; Wang et al., 2022; Haitsiukevich & Ilin, 2022).
@@
Several strategies have been proposed to address these challenges. One approach, introduced by Wight & Zhao (2020), involves time-marching, where the time domain is divided into smaller intervals, with PINNs being sequentially trained on each interval, starting from the initial condition. While effective, this method can be computationally expensive due to the need to train multiple networks. To improve information propagation, researchers have emphasized the importance of weighting initial condition data and residual points near the initial time (Wight & Zhao, 2020; McClenny & Braga-Neto, 2020). McClenny & Braga-Neto (2020) demonstrated that their self-adaptive PINNs can autonomously learn these weights during training, particularly in scenarios with complex initial conditions, such as the wave equation. Further advancements include the “sequence-to-sequence” approach (Krishnapriyan et al., 2021), causality PINN (Wang et al., 2022), artificial viscosity PINN (Coutinho et al., 2023), Co-training PINN (Zhong et al., 2024).
@@
Beyond loss weighting and time marching, structural modifications to PINNs have also been explored. Braga-Neto (2022) introduced Characteristic-Informed Neural Networks (CINNs) for transport equations, where PDE information is integrated into the neural network architecture. A similar approach, albeit for Dirichlet boundary conditions, was proposed in (Lagaris et al., 1998). Our method builds upon these architectural modifications, with a focus on enforcing (near)-exact fit of the initial condition using hard-constrained transformation.
@@ 2 Methodology
We consider the following setup for general time dependent PDEs. Let $u$ be an unknown function defined on $0,T]\times\bar{\Omega}$ with $\Omega\subset\mathbb{R}^{d}$. Here the physical domain $\Omega$ comes with a Lipschitz boundary $\partial\Omega$ and $\bar{\Omega}=\Omega\cup\partial\Omega$. And we denote $\mathbf{x}=\begin{pmatrix}x_{1}&\cdots&x_{d}\end{pmatrix}^{\top}\in\Omega$. Then we say that $u$ is a solution of a time-dependent PDE, if $u$ satisfies the following
$\begin{cases}u_{t}(t,\mathbf{x})+\mathcal{P}[u](t,\mathbf{x})&=f(t,\mathbf{x}),\quad(t,\mathbf{x})\in(0,T]\times\Omega,\\ u(0,\mathbf{x})&=u_{0}(\mathbf{x}),\quad\mathbf{x}\in\bar{\Omega},\\ \mathcal{B}[u](t,\mathbf{x})&=g(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega,\\ \end{cases}.$ (1)
Here the operator $\mathcal{P}$ has order $K$ and it is defined as
$\mathcal{P}[u]=h(u,\{\partial^{k}_{x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{d}^{j_{d}}}u\}_{k=1}^{K}),\quad\text{$h$ is any multivariate function},$
where $\partial^{k}_{x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{d}^{j_{d}}}u$ is a set that contains all of partial derivatives of $u$ with respective to $\mathbf{x}$ of $k^{th}$ order in the multi variate sense, i.e., for the powers of each coordinate, we have
$j_{1},j_{2},\cdots,j_{d}\geq 0\quad\text{and}\quad j_{1}+j_{2}+\cdots+j_{d}=k.$
Furthermore $\mathcal{B}$ is an operator defined for $(t,\mathbf{x})$ on the boundary. We also assume that the functions $u_{0}$, $f$, and $g$ are user inputs that satisfy the desired regularity so that the existence and uniqueness of solutions for such a PDE are guaranteed. We also assume that the compatibility condition, $g(0,\mathbf{x})=u_{0}(\mathbf{x})$ for $\mathbf{x}\in\partial\Omega$, is satisfied to prevent any ill-conditioning of the solution.
@@
$\mathcal{B}[u](t,\mathbf{x};a,b)=au(t,\mathbf{x})+b\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Robin type},$
where $\mathbf{n}$ is the outward normal vector to $\partial\Omega$ and $a,b$ are two fixed constants. Special values of $a$ and $b$ can give two other BC cases, such as when $b=0$ gives $\mathcal{B}[u](t,\mathbf{x})=u(t,\mathbf{x})$ (Dirichlet) and $a=0$ gives the $\mathcal{B}[u](t,\mathbf{x})=\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x})$ (Neumann). The periodic boundary condition gives $d$ different equations, and they are,
$\mathcal{B}[u](t,\mathbf{x})=u(t,\mathbf{x})-u(t,\mathbf{x}+P\mathbf{e}_{i})=0,\quad i=1,\cdots,d.$ (2)
Here $P$ is the period and $\mathbf{e}_{i}$ is the $i^{th}$ standard basis vector in $\mathbb{R}^{d}$, that is,
@@
One can even consider a mixture of two or all of the above BC conditions. For example, we can have $\Omega=\Gamma_{1}\cup\cdots\cup\Gamma_{K}$ with each $\Gamma_{i}$ defines a different kind of BC. In this paper, we will mainly focus on the periodic type of boundary conditions.
Recent advancements in scientific machine learning integrate the principles of Physics, particularly the Physics-derived PDEs, into the training process of machine learning models. This integration enables the development of PINNs to solve for $u$ as follows: find an approximate solution from a set of deep neural networks $\mathcal{H}_{\text{NN}}$, where each network has the same depth, the same number of neurons on each hidden layer, and the same activation functions on each layer, that minimizes the following loss functional
[…]
@@
$\begin{cases}&\text{Data Loss}(u_{nn})=\frac{1}{N_{Data}}\sum_{i=1}^{N_{Data}}|u_{nn}(t^{Data},\mathbf{x}_{i}^{Data})-u(t^{Data},\mathbf{x}_{i}^{Data})|^{2}\\ &\text{IC Loss}(u_{nn})=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2},\\ &\text{BC Loss}(u_{nn})=\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2},\\ &\quad\text{or}\\ &\text{BC Loss}(u_{nn})=\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}|(\mathcal{B}[u_{nn}]-g)(t_{j}^{BC},\mathbf{x}_{j}^{BC})|^{2},\\ &\text{PDE Loss}(u_{nn})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\frac{\partial u}{\partial t}+\mathcal{P}[u_{nn}]-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}\end{cases}$
for $u_{nn}\in\mathcal{H}_{\text{NN}}$. Here $\{(t^{CL},\mathbf{x}^{CL})_{i}\}_{i=1}^{N_{CL}}\in(0,T]\times\Omega$ are called collocation points, $\{(0,\mathbf{x}^{IC})_{i}\}_{i=1}^{N_{IC}}\in\{0\}\times\bar{\Omega}$ are the initial condition points, $\{(t^{BC},\mathbf{x}^{BC})_{i}\}_{i=1}^{N_{BC}}\in[0,T]\times\partial\Omega$ are the boundary condition points, and $\{(t^{Data},\mathbf{x}^{Data})_{i}\}_{i=1}^{N_{Data}}\in(0,T]\times\Omega$ are known data points (they can be noisy). The minimizer, denoted as $u_{NN}$, will an be approximate solution to (1).
2.1 Hard Constraints
The main motivation for us to consider hard constrains transformation for PINNs is the possible ill-conditioning of the PDE solution operator’s dependence on IC and BC. Traditional numerical methods do not encounter this kind of difficulty, as the IC data is exactly satisfied at the starting point of the solution loop. PINNs, on the other hand, use $L^{2}$-loss to fit the IC/BC and PDE residual data, which causes the solutions to be highly sensitive to the tightness of fit for the IC in the training, as pointed out in (McClenny & Braga-Neto, 2020). By transforming the need of fitting IC/BC conditions out of PINNs, it reduces the spectral bias in the training. Therefore, We consider the following transformation
$\tilde{u}(t,\mathbf{x})=\psi(t,\mathbf{x})+\phi(t,\mathbf{x})u_{nn}(t,\mathbf{x}),$ (3)